# Ordinality in Discrete-level Question Difficulty Estimation: Introducing Balanced DRPS and OrderedLogitNN

### v1

@@ Abstract
Keywords Question Difficulty Estimation  $\cdot$ Natural language processing  $\cdot$ Ordinal regression  $\cdot$ Ordered logit  $\cdot$ Fine-tuning
@@ 1 Introduction
Question Difficulty Estimation (QDE), also known as question calibration, aims to predict a question’s difficulty directly from its textual content and its answer options. This task plays a central role in personalized learning tools such as computerized adaptive testing (Van der Linden and Glas, 2000) and dynamic online learning platforms, which aim to present questions aligned with a learner’s proficiency level. Selecting questions that are either too easy or excessively difficult can reduce student motivation and hinder learning outcomes (Wang et al., 2014). Reliable estimation of question difficulty is therefore essential.
Traditionally, QDE has relied on manual calibration (Attali et al., 2014) or pretesting (Lane et al., 2016), both of which are time-consuming and costly. To address these limitations, research has explored the use of natural language processing (NLP) techniques. These approaches train machine learning models to infer difficulty from the question text, allowing for rapid and scalable calibration of new questions without the need for manual intervention.
@@ 2 Related Work
Building on the survey by Benedetto et al. (2023), we further investigate studies in QDE that utilize datasets with discrete difficulty levels. Question difficulty is defined using one of three main approaches: (i) Classical Test Theory (CTT) (Hambleton and Jones, 1993), (ii) Item Response Theory (IRT) (Hambleton, 1991), and (iii) manual calibration. In the case of manual calibration with expert annotators, difficulty is almost exclusively assigned in discrete levels due to its ease of use. While difficulty scores derived from CTT and IRT are continuous by nature, they are often discretized in practical applications to facilitate interpretation. Table 1 provides an overview of related work in discrete-level QDE.
@@
Hsu et al. (2018) 2018 IRT (Adjacent) accuracy
Yang and Suyong (2018) 2018 Manual Accuracy
Fang et al. (2019) 2019 Manual Accuracy
Lin et al. (2019) 2019 Manual, IRT (Adjacent) accuracy
Zhou and Tao (2020) 2020 CTT Accuracy, $F_{1}$-score
Loginova et al. (2021) 2021 Manual Accuracy
Benedetto (2023) 2023 Manual RMSE, $R^{2}$, Spearman’s $\rho$
Thuy et al. (2024) 2024 Manual RMSE
@@ 2.1 Output types
Early work on discrete-level QDE employed classification models, such as support vector machines and Bayesian NNs (Hsu et al., 2018; Yang and Suyong, 2018; Fang et al., 2019). Subsequently, Lin et al. (2019) proposed a discretized regression approach using an LSTM-based NN, where difficulty was predicted as a continuous value between 0 and 1 and mapped to discrete intervals (e.g., $[0.0;0.2)$, $[0.2;0.4)$, etc.).
Zhou and Tao (2020) introduced a multi-task BERT-based model that leverages shared representations across datasets, using a classification head to predict difficulty levels. To reduce the reliance on large labeled datasets in supervised methods, Loginova et al. (2021) proposed an unsupervised QDE method. They leverage the uncertainty in pre-trained question-answering models as a proxy for human-perceived difficulty, computed as the variance over the predictions from an ensemble of classification models.
Benedetto (2023) conducted a benchmarking study comparing traditional machine learning methods and end-to-end NNs across datasets with both discrete and continuous difficulty labels. Their results show that fine-tuned Transformer-based models such as BERT and DistilBERT consistently outperform classical methods. However, these models exclusively employed a discretized regression approach to address the ordinal nature of the labels.
More recently, Thuy et al. (2024) investigated active learning for QDE, demonstrating that comparable performance to fully supervised models can be achieved by labeling only a small fraction of the training data. Yet again, only a discretized regression modeling strategy was considered. As summarized in Table 1, prior work has not directly compared discretized regression and classification approaches, and specialized ordinal regression methods remain entirely unexplored.
@@ 2.2 Metrics
Accuracy is the most widely used evaluation metric for discrete-level QDE, particularly in studies employing classification models, as it aligns with the cross-entropy loss typically used during training. Lin et al. (2019) is the only study using a discretized regression approach that also reports accuracy. However, accuracy fails to account for the ordinal structure of difficulty levels: all misclassifications are treated equally, regardless of their distance from the true label. Even if a prediction is incorrect, it should still be as close as possible to the true difficulty level. Additionally, accuracy is a threshold-based metric, relying solely on the final predicted label rather than the full output distribution—an issue that also applies to metrics such as the $F_{1}$-score. To partially address this, Hsu et al. (2018) and Lin et al. (2019) report adjacent accuracy, defined as the proportion of predictions within $k$ levels of the true label. However, the choice of $k$ is often arbitrary and dataset-dependent, complicating comparisons across studies.
The most recent works, which adopt discretized regression approaches, report RMSE as their primary evaluation metric, treating difficulty levels as integers. While RMSE reflects the ordinal structure to some extent, it assumes uniform spacing between levels—a condition rarely met in real-world data. For example, in primary school, there is a non-linear increase in difficulty over the years (Coe et al., 2008). This limitation also applies to metrics such as $R^{2}$ and Spearman’s rank correlation. Furthermore, since RMSE is used as the loss function in discretized regression models, these models benefit from being evaluated on the same objective they were optimized for, giving them an unfair advantage and introducing a potential bias in comparative evaluations.
In addition to the ordinality aspect, these commonly used metrics fail to account for class imbalance, which is a prevalent issue in discrete-level QDE. In many datasets, mid-range difficulty levels tend to dominate, while questions at the extremes—those that are very easy or very difficult—are underrepresented (Liang et al., 2019; Clark et al., 2018). Standard metrics, which compute aggregate scores across all samples, inherently place greater weight on the majority classes. As a result, model performance on minority classes is often underrepresented, leading to inflated metrics that do not accurately reflect model effectiveness across the full difficulty spectrum. This is particularly problematic in educational contexts, where balanced performance across all difficulty levels is essential to ensure adequate personalized learning experiences.
@@ 3 Balanced Discrete Ranked Probability Score
The Continuous Ranked Probability Score (CRPS) is the most widely adopted scoring rule for evaluating probabilistic forecasts of real-valued variables, such as in precipitation forecasting (Gneiting and Raftery, 2007). It is defined as the integral of the squared difference (i.e., Brier score) between the cumulative distribution function (CDF) of a probabilistic forecast $F$ and the CDF of the observed outcome, at all real-valued thresholds. The observed outcome is represented as a degenerate distribution, as its CDF is a step function. Formally, given a dataset $D=\{\mathbf{x}_{i},y_{i}\}^{N}_{i=1}$, the CRPS is computed as:
@@
This property makes the CRPS particularly well-suited for ordinal prediction tasks. In such cases, the DRPS serves as a natural extension of the CRPS for discrete outcomes across $K$ ordered categories. The DRPS has only been applied in meteorology (Weigel et al., 2007) and has received little attention in the general field of ordinal regression. It is defined as:
@@
In this work, we introduce the balanced DRPS to address the class imbalance in discrete-level QDE datasets, where extreme difficulty levels are typically underrepresented compared to mid-range levels (Liang et al., 2019; Clark et al., 2018). The popular metrics (accuracy and RMSE) and standard DRPS compute unweighted averages, which overemphasize performance on majority classes and can produce misleadingly high scores on imbalanced data. However, for educational practitioners, robust performance across the full spectrum of difficulty levels—including the rarest—is essential. To ensure fair evaluation, the balanced DRPS weights each observation inversely proportional to the prevalence of its true class, $w_{i}=\frac{1}{\sum_{j=1}^{N}\mathds{1}\{y_{j}=y_{i}\}}$, thereby giving equal importance to all difficulty levels. Formally:
@@ 4 Ordered Logit for NNs
OrderedLogitNN extends the classical ordered logit model to NNs, effectively bridging the gap between econometrics and deep learning. At its core, it is a latent variable model, where each observation is associated with an unobserved continuous utility value $y_{i}^{*}$ (Greene and Hensher, 2010) modeled as: $y_{i}^{*}=\mathbf{x}_{i}\mathbf{\beta}+\epsilon_{i}$. The observed ordinal outcome $y_{i}$ is derived from $y_{i}^{*}$ through a censoring mechanism, whereby the continuous latent variable is mapped to one of $K$ discrete categories based on a sequence of $K+1$ increasing threshold values $\{\mu_{-1},\mu_{0},\mu_{1},\dots,\mu_{K-1}\}$.
@@
Figure 3: Ordered logit example.
@@ 5 Existing approaches for ordinal regression with NNs
This section describes existing approaches to handle ordinal regression using NNs. Depending on the specific approach, the amount of ordinal information used and the underlying assumptions vary. This study uses three existing specialized ordinal regression methods that follow the extended binary classification framework, most widely used in the ordinal regression literature (Shi et al., 2023). Note that the methods discussed below are not tied to any specific architecture and can be utilized with any NN.
@@ 5.1 Discretized regression
Using a discretized regression approach in an ordinal QDE problem assumes that the inter-level distances are equal. However, this condition is only rarely satisfied in practice (Coe et al., 2008). On the ARC dataset with levels “grade 3” to “grade 9”, for example, such an approach assumes that the jump in difficulty among all grades is identical.
@@ 5.3 Ordinal: OR-NN
A popular general machine learning approach to ordinal regression is to cast it as an extended binary classification problem (Li and Lin, 2006), leveraging the relative order among the labels. That is, the ordinal regression task with $K$ ranks is represented as a series of $K-1$ simpler binary classification sub-problems. For each rank index $k\in{1,2,\dots,K-1}$, a binary classifier is trained according to whether the rank of a sample is larger than $k$. As such, all $K-1$ tasks share the same intermediate layers but are assigned distinct weight parameters in the output layer. In summary, this framework relies on three steps: (i) extending rank labels to binary vectors, (ii) training binary classifiers on the extended labels, and (iii) computing the predicted rank label from the binary classifiers.
In 2016, the authors of Niu et al. (2016) adapted this framework to train NNs for ordinal regression; we refer to this method as OR-NN. More formally, a rank label $y_{i}$ is first extended into $K-1$ binary vectors $y_{i}^{(1)},\dots,y_{i}^{(K-1)}$ such that the $y_{i}^{(k)}\in\{0,1\}$ indicates whether $y_{i}$ exceeds rank $r_{k}$, for instance, $y_{i}^{(k)}=\mathds{1}\{y_{i}>r_{k}\}$. Using the extended binary labels, a single NN is trained with $K-1$ binary classifiers in the output layer to minimize the cross-entropy loss. Based on the binary task predictions, the predicted rank label is $\hat{y}_{i}=r_{q_{i}}$. The rank index $q_{i}$ is given by
@@
Figure 4: Example of a rank inconsistent and rank consistent prediction on a QDE task with $K=5$ levels. Figure adapted from Shi et al. (2023).
@@ 5.4 Ordinal: CORAL and CORN
CORAL (Cao et al., 2020) achieves rank consistency by imposing a weight-sharing constraint in the last layer. Instead of learning distinct weights between each unit in the penultimate layer and each output unit, CORAL enforces that the $K-1$ binary tasks share the same weight parameters to the units in the penultimate layer. In addition, CORAL learns independent bias terms for each output unit as opposed to a single bias term for the output layer.
@@
However, while CORAL outperforms the OR-NN method in age prediction (Cao et al., 2020), the weight-sharing constraint may restrict the expressiveness and capacity of the NN.
More recently, the authors of Shi et al. (2023) proposed CORN, which guarantees rank consistency without restricting the NN’s expressiveness and capacity. CORN achieves rank consistency by a novel training scheme which uses conditional training sets in order to obtain the unconditional rank probabilities.
@@ 6.1 Data
RACE++ (Liang et al., 2019; Lai et al., 2017) is a dataset of reading comprehension MCQs. Questions are labeled with one of three difficulty levels—1 (middle school), 2 (high school), and 3 (university level)—which we treat as ground truth labels for QDE. The label distribution is imbalanced: 25% of the questions are labeled as level 1, 62% as level 2, and 13% as level 3. The dataset is partitioned into training, validation, and test sets containing $100,568$; $5599$; and $5642$ questions, respectively.
ARC (Clark et al., 2018) is a dataset of science MCQs across grades 3 through 9. Question difficulty is indicated by the target grade level (i.e., 7 levels), which we use as ground truth. The training, validation, and test splits contain $3358$, $862$, and $3530$ questions, respectively. The distribution is highly imbalanced: level 8 appears approximately $1400$ times, level 5 about 700 times, and level 6 only 100 times. To decrease this imbalance, we follow Benedetto (2023) and downsample the two most frequent levels to 500 examples each, resulting in a partially balanced training set of $2293$ questions.
@@ 6.2 Model Architecture
We focus on end-to-end Transformer-based NNs, as they have been shown to outperform traditional NLP approaches that rely on separate feature engineering and modeling stages (Benedetto, 2023). We fine-tune the Transformer BERT (“bert-base-uncased”) on the task of QDE, stacking an output layer on top of the pre-trained language model. During fine-tuning, both the weights of the output head and the pre-trained model are updated. We follow the input encoding of Benedetto (2023) and concatenate the question and the text of all the possible answer choices in a single sentence, divided by separator tokens.
@@
The experiments are implemented in PyTorch (Ansel et al., 2024) using the HuggingFace (Wolf et al., 2020) package, and results are averaged over five independent runs with random seeds.

### v2

# EvalLAC’25: 2nd Workshop on Automatic Evaluation of Learning and Assessment Content, July 26, 2025, Palermo, Italy

@@ Abstract
keywords:
[…]
@@ 1 Introduction
Question Difficulty Estimation (QDE), also known as question calibration, aims to predict a question’s difficulty directly from its textual content and its answer options. This task plays a central role in personalized learning tools such as computerized adaptive testing [1] and dynamic online learning platforms, which aim to present questions aligned with a learner’s proficiency level. Selecting questions that are either too easy or excessively difficult can reduce student motivation and hinder learning outcomes [2]. Reliable estimation of question difficulty is therefore essential.
Traditionally, QDE has relied on manual calibration [3] or pretesting [4], both of which are time-consuming and costly. To address these limitations, research has explored the use of natural language processing (NLP) techniques. These approaches train machine learning models to infer difficulty from the question text, allowing for rapid and scalable calibration of new questions without the need for manual intervention.
@@ 2 Related Work
Building on the survey by [5], we further investigate studies in QDE that utilize datasets with discrete difficulty levels. Question difficulty is defined using one of three main approaches: (i) Classical Test Theory (CTT) [6], (ii) Item Response Theory (IRT) [7], and (iii) manual calibration. In the case of manual calibration with expert annotators, difficulty is almost exclusively assigned in discrete levels due to its ease of use. While difficulty scores derived from CTT and IRT are continuous by nature, they are often discretized in practical applications to facilitate interpretation. Table 1 provides an overview of related work in discrete-level QDE.
@@
[8] 2018 IRT (Adjacent) accuracy
[9] 2018 Manual Accuracy
[10] 2019 Manual Accuracy
[11] 2019 Manual, IRT (Adjacent) accuracy
[12] 2020 CTT Accuracy, $F_{1}$-score
[13] 2021 Manual Accuracy
[14] 2023 Manual RMSE, $R^{2}$, Spearman’s $\rho$
[15] 2024 Manual RMSE
@@ 2.1 Output types
Early work on discrete-level QDE employed classification models, such as support vector machines and Bayesian NNs [8, 9, 10]. Subsequently, [11] proposed a discretized regression approach using an LSTM-based NN, where difficulty was predicted as a continuous value between 0 and 1 and mapped to discrete intervals (e.g., $[0.0;0.2)$, $[0.2;0.4)$, etc.).
[12] introduced a multi-task BERT-based model that leverages shared representations across datasets, using a classification head to predict difficulty levels. To reduce the reliance on large labeled datasets in supervised methods, [13] proposed an unsupervised QDE method. They leverage the uncertainty in pre-trained question-answering models as a proxy for human-perceived difficulty, computed as the variance over the predictions from an ensemble of classification models.
[14] conducted a benchmarking study comparing traditional machine learning methods and end-to-end NNs across datasets with both discrete and continuous difficulty labels. Their results show that fine-tuned Transformer-based models such as BERT and DistilBERT consistently outperform classical methods. However, these models exclusively employed a discretized regression approach to address the ordinal nature of the labels.
More recently, [15] investigated active learning for QDE, demonstrating that comparable performance to fully supervised models can be achieved by labeling only a small fraction of the training data. Yet again, only a discretized regression modeling strategy was considered. As summarized in Table 1, prior work has not directly compared discretized regression and classification approaches, and specialized ordinal regression methods remain entirely unexplored.
@@ 2.2 Metrics
Accuracy is the most widely used evaluation metric for discrete-level QDE, particularly in studies employing classification models, as it aligns with the cross-entropy loss typically used during training. [11] is the only study using a discretized regression approach that also reports accuracy. However, accuracy fails to account for the ordinal structure of difficulty levels: all misclassifications are treated equally, regardless of their distance from the true label. Even if a prediction is incorrect, it should still be as close as possible to the true difficulty level. Additionally, accuracy is a threshold-based metric, relying solely on the final predicted label rather than the full output distribution—an issue that also applies to metrics such as the $F_{1}$-score. To partially address this, [8] and [11] report adjacent accuracy, defined as the proportion of predictions within $k$ levels of the true label. However, the choice of $k$ is often arbitrary and dataset-dependent, complicating comparisons across studies.
The most recent works, which adopt discretized regression approaches, report RMSE as their primary evaluation metric, treating difficulty levels as integers. While RMSE reflects the ordinal structure to some extent, it assumes uniform spacing between levels—a condition rarely met in real-world data. For example, in primary school, there is a non-linear increase in difficulty over the years [16]. This limitation also applies to metrics such as $R^{2}$ and Spearman’s rank correlation. Furthermore, since RMSE is used as the loss function in discretized regression models, these models benefit from being evaluated on the same objective they were optimized for, giving them an unfair advantage and introducing a potential bias in comparative evaluations.
In addition to the ordinality aspect, these commonly used metrics fail to account for class imbalance, which is a prevalent issue in discrete-level QDE. In many datasets, mid-range difficulty levels tend to dominate, while questions at the extremes—those that are very easy or very difficult—are underrepresented [17, 18]. Standard metrics, which compute aggregate scores across all samples, inherently place greater weight on the majority classes. As a result, model performance on minority classes is often underrepresented, leading to inflated metrics that do not accurately reflect model effectiveness across the full difficulty spectrum. This is particularly problematic in educational contexts, where balanced performance across all difficulty levels is essential to ensure adequate personalized learning experiences.
@@ 3 Balanced Discrete Ranked Probability Score
The Continuous Ranked Probability Score (CRPS) is the most widely adopted scoring rule for evaluating probabilistic forecasts of real-valued variables, such as in precipitation forecasting [19]. It is defined as the integral of the squared difference (i.e., Brier score) between the cumulative distribution function (CDF) of a probabilistic forecast $F$ and the CDF of the observed outcome, at all real-valued thresholds. The observed outcome is represented as a degenerate distribution, as its CDF is a step function. Formally, given a dataset $D=\{\mathbf{x}_{i},y_{i}\}^{N}_{i=1}$, the CRPS is computed as:
@@
This property makes the CRPS particularly well-suited for ordinal prediction tasks. In such cases, the DRPS serves as a natural extension of the CRPS for discrete outcomes across $K$ ordered categories. The DRPS has only been applied in meteorology [20] and has received little attention in the general field of ordinal regression. It is defined as:
@@
In this work, we introduce the balanced DRPS to address the class imbalance in discrete-level QDE datasets, where extreme difficulty levels are typically underrepresented compared to mid-range levels [17, 18]. The popular metrics (accuracy and RMSE) and standard DRPS compute unweighted averages, which overemphasize performance on majority classes and can produce misleadingly high scores on imbalanced data. However, for educational practitioners, robust performance across the full spectrum of difficulty levels—including the rarest—is essential. To ensure fair evaluation, the balanced DRPS weights each observation inversely proportional to the prevalence of its true class, $w_{i}=\frac{1}{\sum_{j=1}^{N}\mathds{1}\{y_{j}=y_{i}\}}$, thereby giving equal importance to all difficulty levels. Formally:
@@ 4 Ordered Logit for NNs
OrderedLogitNN extends the classical ordered logit model to NNs, effectively bridging the gap between econometrics and deep learning. At its core, it is a latent variable model, where each observation is associated with an unobserved continuous utility value $y_{i}^{*}$ [21] modeled as: $y_{i}^{*}=\mathbf{x}_{i}\mathbf{\beta}+\epsilon_{i}$. The observed ordinal outcome $y_{i}$ is derived from $y_{i}^{*}$ through a censoring mechanism, whereby the continuous latent variable is mapped to one of $K$ discrete categories based on a sequence of $K+1$ increasing threshold values $\{\mu_{-1},\mu_{0},\mu_{1},\dots,\mu_{K-1}\}$.
@@
Figure 3: Ordered logit example for three ordinal categories.
@@ 5 Existing approaches for ordinal regression with NNs
This section describes existing approaches to handle ordinal regression using NNs. Depending on the specific approach, the amount of ordinal information used and the underlying assumptions vary. This study uses three existing specialized ordinal regression methods that follow the extended binary classification framework, most widely used in the ordinal regression literature [22]. Note that the methods discussed below are not tied to any specific architecture and can be utilized with any NN.
@@ 5.1 Discretized regression
Using a discretized regression approach in an ordinal QDE problem assumes that the inter-level distances are equal. However, this condition is only rarely satisfied in practice [16]. On the ARC dataset with levels “grade 3” to “grade 9”, for example, such an approach assumes that the jump in difficulty among all grades is identical.
@@ 5.3 Ordinal: OR-NN
A popular general machine learning approach to ordinal regression is to cast it as an extended binary classification problem [23], leveraging the relative order among the labels. That is, the ordinal regression task with $K$ ranks is represented as a series of $K-1$ simpler binary classification sub-problems. For each rank index $k\in{1,2,\dots,K-1}$, a binary classifier is trained according to whether the rank of a sample is larger than $k$. As such, all $K-1$ tasks share the same intermediate layers but are assigned distinct weight parameters in the output layer. In summary, this framework relies on three steps: (i) extending rank labels to binary vectors, (ii) training binary classifiers on the extended labels, and (iii) computing the predicted rank label from the binary classifiers.
In 2016, the authors of [24] adapted this framework to train NNs for ordinal regression; we refer to this method as OR-NN. More formally, a rank label $y_{i}$ is first extended into $K-1$ binary vectors $y_{i}^{(1)},\dots,y_{i}^{(K-1)}$ such that the $y_{i}^{(k)}\in\{0,1\}$ indicates whether $y_{i}$ exceeds rank $r_{k}$, for instance, $y_{i}^{(k)}=\mathds{1}\{y_{i}>r_{k}\}$. Using the extended binary labels, a single NN is trained with $K-1$ binary classifiers in the output layer to minimize the cross-entropy loss. Based on the binary task predictions, the predicted rank label is $\hat{y}_{i}=r_{q_{i}}$. The rank index $q_{i}$ is given by
@@
Figure 4: Example of a rank inconsistent and rank consistent prediction on a QDE task with $K=5$ levels. Figure adapted from [22].
@@ 5.4 Ordinal: CORAL and CORN
CORAL [25] achieves rank consistency by imposing a weight-sharing constraint in the last layer. Instead of learning distinct weights between each unit in the penultimate layer and each output unit, CORAL enforces that the $K-1$ binary tasks share the same weight parameters to the units in the penultimate layer. In addition, CORAL learns independent bias terms for each output unit as opposed to a single bias term for the output layer.
@@
However, while CORAL outperforms the OR-NN method in age prediction [25], the weight-sharing constraint may restrict the expressiveness and capacity of the NN.
More recently, the authors of [22] proposed CORN, which guarantees rank consistency without restricting the NN’s expressiveness and capacity. CORN achieves rank consistency by a novel training scheme which uses conditional training sets in order to obtain the unconditional rank probabilities.
@@ 6.1 Data
RACE++ [17, 26] is a dataset of reading comprehension MCQs. Questions are labeled with one of three difficulty levels—1 (middle school), 2 (high school), and 3 (university level)—which we treat as ground truth labels for QDE. The label distribution is imbalanced: 25% of the questions are labeled as level 1, 62% as level 2, and 13% as level 3. The dataset is partitioned into training, validation, and test sets containing $100,568$; $5599$; and $5642$ questions, respectively.
ARC [18] is a dataset of science MCQs across grades 3 through 9. Question difficulty is indicated by the target grade level (i.e., 7 levels), which we use as ground truth. The training, validation, and test splits contain $3358$, $862$, and $3530$ questions, respectively. The distribution is highly imbalanced: level 8 appears approximately $1400$ times, level 5 about 700 times, and level 6 only 100 times. To decrease this imbalance, we follow [14] and downsample the two most frequent levels to 500 examples each, resulting in a partially balanced training set of $2293$ questions.
@@ 6.2 Model Architecture
We focus on end-to-end Transformer-based NNs, as they have been shown to outperform traditional NLP approaches that rely on separate feature engineering and modeling stages [14]. We fine-tune the Transformer BERT (“bert-base-uncased”) on the task of QDE, stacking an output layer on top of the pre-trained language model. During fine-tuning, both the weights of the output head and the pre-trained model are updated. We follow the input encoding of [14] and concatenate the question and the text of all the possible answer choices in a single sentence, divided by separator tokens.
@@
The experiments are implemented in PyTorch [27] using the HuggingFace [28] package, and results are averaged over five independent runs with random seeds.